Dundalk Industries operates production facilities in ;several locations across maryland. Average monthly output and cost data for three facilities are shown in the table below.

A. calculate the labor productivity of each facility

b. for each facility , calculate the multifactor productivity for labor and materials

c. if the company needed to close on of the plants, which of the three facilities should be closed

location A B C
finished goods (units) 20,000 15,000 25,000
labor cost 3,500 4,200 2,500
material costs 3,500 3,000 2000
energy costs 1,000 1,500 1,200
transportation 250 2500 2000
overhead 1200 3000 2500
Answer #1

For Location A -

Output = 20000 units

Labor Input = 3500

Overall Input = 3500 + 3500 + 1000 + 250 + 1200 = 9450

Productivity = Input / Output

Labor Productivity = 20000/3500 = 5.71 unit per $

Multifactor Productivity = 20000/9450 = 2.12 unit per $

For Location B -

Output = 15000 units

Labor Input = 4200

Overall Input = 4200 + 3500 + 1500 + 2500 + 3000 = 14700

Productivity = Input / Output

Labor Productivity = 15000/4200 = 3.57 unit per $

Multifactor Productivity = 15000/14700 = 1.02 unit per $

For Location C -

Output = 25000 units

Labor Input = 2500

Overall Input = 2500 + 2000 + 1200 + 2000 + 2500 = 10200

Productivity = Input / Output

Labor Productivity = 25000/2500 = 10 unit per $

Multifactor Productivity = 25000/10200 = 2.45 unit per $

Since Productivity is least for Location B hence Location B should be closed.